OpenStudy (nixonn):

What are the intercepts of y = 4x − 2?

OpenStudy (anonymous):

your mom

OpenStudy (nixonn):

omfg mddie

OpenStudy (anonymous):

haha js

OpenStudy (anonymous):

y intercept set x=0 y=4(0)-2 y=-2 y intercept is (0,-2)

OpenStudy (anonymous):

now solve for the y-intercept y=0

OpenStudy (anonymous):

wait solve for x intercept jaja my bad y=0 4x-2=0 add 2 to both sides 4x=2 divide by 4 x=2/4 simplify x=1/2

OpenStudy (anonymous):

if that helps

OpenStudy (anonymous):

yeah what they said ^